Solve the system
Y = -5x
X +0.8*y = -7.2

Found 3 solutions by checkley79, Dionysus10474, solver91311:
Answer by checkley79(3341)

Y = -5x
X +0.8y = -7.2 REPLACE Y WITH -5X IN THE SECOND EQUATION & SOLVE FOR X.
X+.8(-5X)=-7.2
X-4X=-7.2
-3X=-7.2
X=-7.2/-3
X=2.4 ANS.
Y=-5*2.4
Y=-12 ANS.
PROOF:
2.4+.8*-12=-7.2
2.4-9.6=-7.2
-7.2=-7.2

Answer by Dionysus10474(5)

y = -5x
x + 0.8*y = -7.2
To solve this system you need to put both equations in one equation.
First equation: y = -5x
so 2*y = 2*(-5x)
and therefore 0.8*y = 0.8*(-5x)
which we have in the second equation, so we can put 0.8*y in the second equation.
x + (0.8*(-5x)) = -7.2
x + -4x = -7.2
-3x = -7.2
x = -7.2/-3
x = 2.4
Put x in the first equation and you get the value of y.
y = -5x
y = -5*2.4
y = -12

Answer by solver91311(24713)



You can't solve a system with 4 different variables and only 2 equations.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism